Assume that in 2018, the first edition of a comic book was sold at auction for $797,800. The comic book was originally sold in 1939 for $.10. For this to have been true, what was the annual increase in the value of the comic book? (Do not round intermediate calculations and enter your answer as a percent rounded to 2 decimal places, e.g., 32.16.)
Answer #1

We use the formula:
A=P(1+r/100)^n
where
A=future value
P=present value
r=rate of interest
n=time period.

797,800=0.1*(1+r/100)^79

(797,800/0.1)^(1/79)=1+r/100

1+r/100=1.2228

r=1.2228-1

=22.28%(Approx).